40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Sat Mar 12 05:32:40 2011, Zach, Laser, GYRO, current gyro noise gyro_noise_with_contribs_3_11_11.pngoutput_box.pnginput_box.png
    Reply  Sat Mar 12 16:11:35 2011, Koji, Laser, GYRO, current gyro noise 
       Reply  Sat Mar 12 16:22:53 2011, Zach, Laser, GYRO, current gyro noise 
          Reply  Sat Mar 12 19:20:19 2011, Koji, Laser, GYRO, current gyro noise 
             Reply  Sat Mar 12 21:24:14 2011, Zach, Laser, GYRO, current gyro noise OLTFs_3_11_11.png
                Reply  Sun Mar 13 00:27:29 2011, Koji, Laser, GYRO, current gyro noise 
Message ID: 1350     Entry time: Sat Mar 12 21:24:14 2011     In reply to: 1349     Reply to this: 1351
Author: Zach 
Type: Laser 
Category: GYRO 
Subject: current gyro noise 

I'm not sure I understand your points..

"- This AOM feedback should agree with the calibrated error signal of the secondary cavity.  This is not confirmed."

I think you mean that the secondary error signal calibrated into Hz via the optical response ([V/Hz]) should be the same as the suppressed noise calculated by multiplying the AOM feedback signal through the AOM response ([Hz/V]) and dividing by the OLTF, as I did for the primary loop here. If so, then no, I haven't done this yet. I was about to do it now but I must have left the CCW error signal disconnected from the DAQ 

 

"- This noise is suppressed by the loop gain of the secondary cavity. (The spillover noise of the secondary.) This is not shown. Also the OLTF of the secondary is not shown."

I don't think I know what you mean here. The spillover noise from the primary cavity (which can be calculated from the primary error signal) is present in the AOM actuation signal, as it tries to cancel the noise that the primary loop has failed to.

Here are both of the OLTFs:

OLTFs_3_11_11.png

 

"- The AOM feedback at around 100Hz shows the same level as the primary spillover. Does this suggest that the loop gain of the secondary is higher than 1k???"

In frequency bands where the residual common-mode noise is higher than the differential noise, the AOM feedback spectrum should be the same as the calculated spillover noise (blue and red, respectively). This appears to be true above ~100 Hz. It was true above 30 Hz in the old budget. It makes sense that the residual common-mode noise has gone down a little below 100 Hz, because the primary OL gain has gone up a bit from the extra integrator I added, but it is strange that there is still about the same level of noise below 100 Hz in the AOM signal as there was before, despite the fact that we replaced PDs. It could be that the optical gain is still low enough that electronics noise dominates here, but we won't know that until I plot the PD/demod noise alongside it. I will do that tonight or tomorrow.

"- Of course we like to see the OLTF of the PLL."

I plan to measure this, but it is a little difficult since I don't have an electrical summation point for this as I do for the locking loops. I think I can do it by sweeping the AOM frequency (with the secondary loop open, that is), but I haven't worked it out yet.

 

Quote:

I understand your explanations. Now I feel these points.

- This AOM feedback should agree with the calibrated error signal of the secondary cavity.  This is not confirmed.

- This noise is suppressed by the loop gain of the secondary cavity. (The spillover noise of the secondary.) This is not shown. Also the OLTF of the secondary is not shown.

- The AOM feedback at around 100Hz shows the same level as the primary spillover. Does this suggest that the loop gain of the secondary is higher than 1k???

- Of course we like to see the OLTF of the PLL.

Are you working with the measurements, Zach?

 

 

 

ELOG V3.1.3-